# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v1

@@ Abstract
We consider channels with synchronization errors modeled as insertions and deletions. A classical result for such channels is the information stability of such channels, hence the existence of the Shannon capacity, when the synchronization errors are memoryless. In this paper, we extend this result to the case where the insertions and deletions have memory. Specifically, we assume that the synchronization errors are governed by a stationary and ergodic finite state Markov chain, and prove that mutual information capacity of such channels exist, and it is equal to its coding capacity, showing that there exists a coding scheme which achieves this limit.
1 Introduction
@@
A mathematical model of a physical channel is defined by conditional distributions $W(y|x)$ for all $x\in{\mathcal{X}},y\in{\mathcal{Y}}$, which means that if a symbol $x\in{\mathcal{X}}$ is input, output $y\in{\mathcal{Y}}$ is observed with probability [^1] This is valid only for a countable ${\mathcal{Y}}$; for uncountable case, there should be output PDFs instead; in this paper uncountable ${\mathcal{Y}}$ are not considered. $W(y|x)$. The fact that $W$ is a channel with the input set ${\mathcal{X}}$ and the output set ${\mathcal{Y}}$ will be denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$. The famous Shannon’s theorem [1] considers a memoryless channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ that is used $n$ times independently. The equivalent channel that transmits $n$ symbols simultaneously is the channel $W^{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}^{n}$, where the conditional probabilities are given by $W^{n}(y_{1}^{n}|x_{1}^{n})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, where we use vector notation $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$.
@@
Consider now an insertion channel $V_{n}$ with $n$ input symbols $x\in{\mathcal{X}}^{n}$. In general, for each input symbol $x_{i}\in{\mathcal{X}}$, an output vector $y_{i}$ over ${\mathcal{X}}$ of length $\geq 1$ is produced. In contrast with the channel $W^{n}$, the resulting output of $V_{n}$ is a concatenation of output vectors, but not the vector $y_{1}^{n}$ of output vectors. For example, if input bits are $(1,0,0)$ and the corresponding outputs for each bit are $(1,0)$, $(0)$ and $(1,0)$, then the overall channel output would be $(1,0,0,1,0)$, not $((1,0),(0),(1,0))$.
@@
In general, a communication setup, besides conditional probabilities of the original channel, also includes some underlying method of combining the outputs of multiple channel uses. In the case of simple memoryless channels, the combining method is just stacking together output symbols into the output vector; in the case of insertion channel, the combining method is concatenation.
In general we can say that the transition probabilities for multiple uses of a channel are not uniquely defined by the conditional probabilities of the single-use (or “original”) channel. For this reason, in the non-trivial setup, instead of a channel, a channel sequence $(W_{n}:{\mathcal{X}}_{n}\rightsquigarrow{\mathcal{Y}}_{n})_{n\in{\mathbb{N}}}$ is usually considered, where sets ${\mathcal{X}}_{n}$, ${\mathcal{Y}}_{n}$ can vary for different $n$. For simplicity, we consider the case when input set for $W_{n}$ is ${\mathcal{X}}_{n}={\mathcal{X}}^{n}$ and ${\mathcal{X}}$ is fixed for all $n$.
@@
For a channel sequence the information capacity can be defined as $\lim_{n\to\infty}{\mathbf{I}}(W_{n})/n$, where ${\mathbf{I}}(W_{n})$ denotes the maximum mutual information between the input and the output of a channel $W_{n}$ over all possible input distributions. The coding capacity (or operational capacity) is the (maximum) asymptotic rate of coding schemes which achieve arbitrary small error probability as $n\to\infty$. The capacity theorem is a statement that the information capacity is equal to coding capacity.
@@
There are many capacity theorems that generalize the Shannon’s theorem on non-trivial channel sequences. Dobrushin’s work [2] considers the most general case, where input and output alphabets can be continuous, and the metric of closeness of decoder’s output to the original signal can be arbitrary . It is shown that the so-called “information stability” of a channel sequence is sufficient for the capacity theorem to hold. In [3], Dobrushin proves the capacity theorem for channels with independent synchronization errors. In that model, for each input symbol, an output vector of finite expected length is produced, and all the output vectors are concatenated to form the resulting output. Since this channel model is general and includes any insertion/deletion/substitution errors, we call such model an IDS channel. The generalization of the capacity theorem to the case of IDS channels with continuous input alphabet is studied in [4]. Capacity theorems are also proven for finite-state Markov channels (FSMC) [5], for Gilbert-Elliott channels [6], for deletion channel concatenated with FSMC [7].
@@
IDS channels are used to model different communication channels, including bit patterned media recording [8] and DNA storage channels [9]. For the case of memoryless insertion/deletion channels, capacity upper bounds and lower bounds are also obtained in [10, 11, 12], and the asymptotic expansion of the i.i.d. deletion channel capacity is found in [13]. See [14, 15] for a detailed review of progress in this area. While most existing results focus on the scenario with independent synchronization errors, it is also important to study the case with insertions/deletions with memory. For instance, current reading and writing technologies in DNA sequencing introduce memory to the channel errors [16, 17]. A simple and useful model for synchronization errors with memory is a Markov chain.
@@
In this paper, we consider IDS channels where the underlying insertions/deletions are governed by a Markov chain, referred to as a Markov-IDS channel. That is, each state of the Markov chain is an IDS channel, called component channel, and the overall output of Markov-IDS channel is obtained by concatenating the individual component channel outputs. The Markov-IDS channel is a generalization of both IDS and FSMC: it is more general than IDS since it changes over time according to Markov chain rule; it is more general then FSMC since it concatenates the outputs of each component IDS channel (which can be translated to non-concatenated version via inserting a special symbol at the end of each component channel’s output). We note that while this model generalizes the synchronization errors as studied by Dobrushin, it is not sufficiently general to encompass some DNA channel models with insertions/deletions with memory as in [16, 17] since the channel states do not depend on the channel input in our model. Further extensions are left for future work.
@@
We also introduce some unified notation, which helps to elucidate the idea behind Dobrushin’s methods, which can be seen as applying functions to channel sequences. Namely, if a function is applied on a channel’s output, and this function is not “doing much” in a certain sense, then it changes neither the mutual information capacity nor the coding capacity of the original channel. This means that one can apply any finite chain of such functions and still have the same capacities.
@@
The paper is organized as follows. In Section 2, notations and information-theoretic definitions are introduced. In Section 3, functions on channels and their properties are described. In Section 4, the Markov-IDS channel sequence is defined. In Section 5, the capacity theorem for Markov-IDS channel sequence is proven. In Section 6, the conclusion is drawn.
2 Background
2.1 Notations
@@
We use the following notations. Sets are written in calligraphic letters ${\mathcal{X}},{\mathcal{Y}},{\overline{\mathcal{Y}}}$; operators are written in bold letters ${\mathbf{C}},{\mathbf{E}},{\mathbf{I}}$. For $b,c\in{\mathbb{N}}$, we denote a vector $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$. Here, elements $a_{i}$ can be of any type: a number, a distribution, a channel, etc. Also, we denote infinite sequences as $a^{\infty}=(a_{1},a_{2},\dots)$. For a set ${\mathcal{A}}$, we denote a set of all finite sequences with elements from ${\mathcal{A}}$ by ${\overline{\mathcal{A}}}=\left\{{()}\right\}\cup\bigcup_{i=1}^{\infty}{\mathcal{A}}^{i}.$ For a vector $a=a_{1}^{n}\in{\overline{\mathcal{A}}}$ we denote by ${\mathbf{N}}(a)=n$ the vector length. We denote the concatenation (gluing) operator $\mathbf{g}$ as $\mathbf{g}(a_{1}^{n},b_{1}^{m})=(a_{1},a_{2},...,a_{n},b_{1},...,b_{m}),$ which is also generalized onto any number of input vectors. Also, define the function $\mathbf{g}_{n}:{\overline{\mathcal{A}}}^{n}\to{\overline{\mathcal{A}}}$, which glues $n$ vectors ${\overline{a}}_{i}\in{\overline{\mathcal{A}}}$ as $\mathbf{g}_{n}({\overline{a}}_{1}^{n})=\mathbf{g}({\overline{a}}_{1},{\overline{a}}_{2},...,{\overline{a}}_{n}).$ The inverse image $\mathbf{g}^{-1}_{n}(a)$ is the set of all possible ways of cutting $a\in{\overline{\mathcal{A}}}$ into $n$ subvectors, i.e., $\mathbf{g}^{-1}_{n}(a)=\left\{{{\overline{a}}_{1}^{n}|\mathbf{g}_{n}({\overline{a}}_{1}^{n})=a}\right\}$. For a set ${\mathcal{S}}$, we denote by ${\mathbb{D}}_{\mathcal{S}}$ the set of all distributions over ${\mathcal{S}}$, i.e.
@@
The probability distribution for any probability, expectation or entropy is written as a subscript. We assume summation of all values of random variables, which are not fixed. For example, if we have a conditional distribution $W(u,x|v,y)$, where $u\in{\mathcal{U}},v\in{\mathcal{V}},x\in{\mathcal{X}},y\in{\mathcal{Y}}$, and some specific values of $u$ and $v$ (or fixed them by $\exists$,$\forall$ , ∀ symbols). Then,
@@
Entropy of $x$ is expectation of $-\log_{2}x$, so ${\mathbf{H}}_{A}{\left[{B}\right]}={\mathbf{E}}_{A}{\left[{-\log_{2}B}\right]}$. If $A=B$, then we omit the subscript for simplicity:
@@
2.2 Channels and channel sequences
@@
A channel sequence $W$ is defined as an infinite sequence of channels
$W=(W_{n})_{n\in{\mathbb{N}}},\;W_{i}:{\mathcal{X}}^{i}\rightsquigarrow{\mathcal{Y}}_{i}\;\forall i\in{\mathbb{N}}.$
Note that for any $i$, the output set ${\mathcal{Y}}_{i}$ is arbitrary, but the input set for the $i$-th channel is ${\mathcal{X}}^{i}$, so the $i$-th channel transmits exactly $i$ symbols from ${\mathcal{X}}$.
@@
2.3 Mutual information capacity
The mutual information density of channel $W$ for input distribution $p\in{\mathbb{D}}_{\mathcal{X}}$ and given values of input $x\in{\mathcal{X}}$ and output $y\in{\mathcal{Y}}$ is defined as:
@@ 2.3 Mutual information capacity
The mutual information capacity/i-capacity [2] of channel $W$ is the maximum possible mutual information of $W$ under any input distribution, and is given by [^2] In [2], the setup is for the much more general case, and “$\sup$” over the input distributions is used instead of “$\max$”. In our setup we assumed (see Section 2.2) that ${\mathcal{X}}$ is finite and ${\mathcal{Y}}$ is countable, so the maximum is always achieved as a maximum of a continuous function over a compact set.
@@ 2.4 Coding capacity
This definition obviously implies that if there exists a $(W,M,\varepsilon)$ coding scheme, there also exists a $(W,M^{\prime},\varepsilon^{\prime})$-coding scheme for any $M^{\prime}\leq M$ and $\varepsilon^{\prime}\geq\varepsilon$.
Consider a channel sequence $W$, where $W_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$. We call $R\in{\mathbb{R}}_{\geq 0}$ an achievable rate over $W$, if
@@
We denote by ${\mathbf{R}}(W)$ the set of achievable rates $R$ for channel sequence $W$. By the coding capacity/c-capacity ${\mathbf{C}}(W)$ of the channel sequence $W$ we mean the supremum of all achievable rates [3]:
${\mathbf{C}}(W)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\sup{\mathbf{R}}(W)$ (8)
Note that, $0\in{\mathbf{R}}(W)$ for any $W$, and if $R\in{\mathbf{R}}(W)$, then $R^{\prime}\in{\mathbf{R}}(W)$ for any rate $R^{\prime}<R$. Thus, the set ${\mathbf{R}}(W)$ is a connected set, i.e., there always exists $\alpha\in[0,\log_{2}|{\mathcal{X}}|]$, such that ${\mathbf{R}}(W)\in\left\{{[0,\alpha],[0,\alpha)}\right\}$. This implies that ${\mathbf{C}}(W)=\alpha$, i.e., any channel sequence has the c-capacity.
In [18], the coding capacity (c-capacity) ${\mathbf{C}}(W)$ is derived in terms of mutual information density:
${\mathbf{C}}(W)=\sup\left\{{\alpha\left|\lim_{n\to\infty}~{}\min_{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\Pr_{p_{n}\circ W_{n}(x,y)}\left\{{\frac{\mathbf{i}(W_{n},p_{n},x,y)}{n}\leq\alpha}\right\}=0\right.}\right\}$ (9)
2.5 Information-stable channels
The channel sequence $W$ is called information-stable, if it has i-capacity, i.e., ${\mathbf{I}}(W)$ exists, and for optimal distributions $p^{*}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$ for $W_{n}$, for any $\delta>0$:
$\lim_{n\to\infty}~{}\Pr_{p^{*}_{n}\circ W_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(W_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(W)}\right|>\delta}\right]}=0.$ (10)
Proposition 1 (Theorem 1 in [2]).
If a channel sequence $W$ is information-stable, then ${\mathbf{C}}(W)={\mathbf{I}}(W)$.
3 Functions of channels and their properties
For a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and function $f:{\mathcal{Y}}\to{\mathcal{Z}}$ we write $W^{\prime}={f}{\left({W}\right)}$ to refer to a channel $W^{\prime}:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, defined as a composition of channel $W$ and function $f$. That is,

### v3

@@ Abstract
We consider channels with synchronization errors modeled as insertions and deletions. A classical result for such channels is their information stability, hence the existence of the Shannon capacity, when the synchronization errors are memoryless. In this paper, we extend this result to the case where the insertions and deletions have memory. Specifically, we assume that the synchronization errors are governed by a stationary and ergodic finite state Markov chain, and prove that such channel is information-stable, which implies the existence of a coding scheme which achieves the limit of mutual information. This result implies the existence of the Shannon capacity for a wide range of channels with synchronization errors, with different applications including DNA storage. The methods developed may also be useful to prove other coding theorems for non-trivial channel sequences.
I Introduction
@@
A mathematical model of a physical channel is defined by conditional distributions $W(y|x)$ for all $x\in{\mathcal{X}},y\in{\mathcal{Y}}$, which means that if a symbol $x\in{\mathcal{X}}$ is input, output $y\in{\mathcal{Y}}$ is observed with probability [^2] This is valid only for a countable ${\mathcal{Y}}$; for uncountable case, there should be output PDFs instead; in this paper uncountable ${\mathcal{Y}}$ are not considered. $W(y|x)$. The fact that $W$ is a channel with the input set ${\mathcal{X}}$ and the output set ${\mathcal{Y}}$ will be denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$. Shannon’s theorem [1] considers a memoryless channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ that is used $n$ times independently. The equivalent channel that transmits $n$ symbols simultaneously is the channel $W^{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}^{n}$, where the conditional probabilities are given by $W^{n}(y_{1}^{n}|x_{1}^{n})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, with the use of the vector notation $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$.
@@
In this paper, our focus is on channels with synchronization errors, modeled as insertions, deletions and channel errors. As an example, consider an insertion channel $V_{n}$ with $n$ input symbols $x\in{\mathcal{X}}^{n}$. In general, for each input symbol $x_{i}\in{\mathcal{X}}$, an output vector $y_{i}$ over ${\mathcal{X}}$ of length $\geq 1$ is produced. In contrast with the channel $W^{n}$, the resulting output of $V_{n}$ is a concatenation of output vectors, but not the vector $y_{1}^{n}$ of output vectors. For example, if input bits are $(1,0,0)$ and the corresponding outputs for each bit are $(1,0)$, $(0)$ and $(1,0)$, then the overall channel output would be $(1,0,0,1,0)$, not $((1,0),(0),(1,0))$.
@@
In general, a communication setup, besides conditional probabilities of the original channel, also includes some underlying method of combining the outputs of multiple channel uses. In the case of simple memoryless channels, the combining method is just stacking together output symbols into the output vector; in the case of insertion (and/or deletion) channel, the combining method is concatenation. For this reason, in the non-trivial setup a general case of a channel sequence $(W_{n}:{\mathcal{X}}_{n}\rightsquigarrow{\mathcal{Y}}_{n})_{n\in{\mathbb{N}}}$ is usually considered [2, 3], where sets ${\mathcal{X}}_{n}$, ${\mathcal{Y}}_{n}$ can vary for different $n$. For simplicity, we consider the case when input set for $W_{n}$ is ${\mathcal{X}}_{n}={\mathcal{X}}^{n}$ and ${\mathcal{X}}$ is fixed for all $n$.
@@
For a channel sequence the mutual information capacity can be defined as $\lim_{n\to\infty}{\mathbf{I}}(W_{n})/n$, where ${\mathbf{I}}(W_{n})$ denotes the maximum mutual information between the input and the output of a channel $W_{n}$ over all possible input distributions. The coding capacity (or operational capacity) is the (maximum) asymptotic rate of coding schemes which achieve arbitrary small error probability as $n\to\infty$. The capacity theorem is a statement that the mutual information capacity is equal to the coding capacity.
@@
There are many capacity theorems that generalize the Shannon’s theorem on non-trivial channel sequences. Dobrushin’s work [4] considers the most general case, where input and output alphabets can be continuous, and the metric of closeness of decoder’s output to the original signal can be arbitrary. It is shown that the so-called “information stability” of a channel sequence is sufficient for the capacity theorem to hold. Later, in [2] it was shown that information stability is necessary and sufficient for the capacity theorem. In [5], Dobrushin proves the capacity theorem for channels with independent synchronization errors. In that model, for each input symbol, an output vector of finite expected length is produced, and all the output vectors are concatenated to form the resulting output. Since this channel model includes any insertion/deletion/substitution errors, we call such model an IDS channel. The generalization of the capacity theorem to the case of IDS channels with continuous input alphabet is studied in [6]. Capacity theorems are also proven for finite-state Markov channels (FSMC) [7], for Gilbert-Elliott channels [8], for deletion channels concatenated with FSMC [9], for IDS channels with intersymbol interference [10].
@@
IDS channels are used to model various communication channels, including bit patterned media recording [11] and DNA storage channels [12]. For the case of memoryless insertion/deletion channels, capacity upper bounds and lower bounds are also obtained in [13, 14, 15], and the asymptotic expansion of the independent identically distributed deletion channel capacity is found in [16]. See [17, 18] for a detailed review of progress in this area. While most existing results focus on the scenario with independent synchronization errors, it is also important to study the case with insertions/deletions with memory. For instance, current reading and writing technologies in DNA sequencing introduce memory to the channel errors [19, 20].
@@
In this paper, we consider IDS channels with Markov memory, i.e., the underlying insertions/deletions are governed by a Markov chain. This model is referred to as a Markov-IDS channel. That is, each state of the Markov chain is an IDS channel, called component channel, and the overall output of Markov-IDS channel is obtained by concatenating the individual component channel outputs. The Markov-IDS channel is a generalization of both IDS and FSMC: it is more general than IDS since statistics of the synchronization errors change over time according to Markov chain rule; it is more general then FSMC since it concatenates the outputs of each component IDS channel (which can be translated to non-concatenated version via inserting a special symbol at the end of each component channel’s output). We note that while this model generalizes the synchronization errors as studied by Dobrushin, it is not sufficiently general to encompass some DNA channel models with insertions/deletions with memory as in [19, 20] since the channel states do not depend on the channel input in our model. Alternatively, in [10], the capacity theorem is proven for IDS channels with intersymbol interference (ISI). The difference between the ISI-IDS channel model in [10] and Markov-IDS channel model in this paper is that the state of ISI-IDS channel is defined by a few recent input bits, and the channel state in the Markov-IDS model is random (though the distribution is defined by the previous state).
@@
In this paper, we also introduce a unified notation, which helps to elucidate the idea behind Dobrushin’s methods, which can be seen as applying functions to channel sequences. Namely, if a function is applied on a channel’s output, and this function is not “doing much” in a certain sense, then the modified channel has the same mutual information capacity, coding capacity and information stability, as the original channel. This means that one can apply any finite number of such functions and still have the same capacities.
@@
The paper is organized as follows. In Section II, notations and information-theoretic definitions are introduced. In Section III, functions on channels and their properties are described. In Section IV, the Markov-IDS channel sequence is defined. In Section V, the capacity theorem for Markov-IDS channel sequence is proven. In Section VI, conclusions are drawn.
II Background
II-A Notations
@@
We use the following notation. Sets are written in calligraphic letters ${\mathcal{X}},{\mathcal{Y}},{\overline{\mathcal{Y}}}$; operators are written in bold letters ${\mathbf{C}},{\mathbf{E}},{\mathbf{I}}$. For $b,c\in{\mathbb{N}}$, we denote a vector $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$. Here, elements $a_{i}$ can be of any type: a number, a distribution, a channel, etc. Also, we denote infinite sequences as $a^{\infty}=(a_{1},a_{2},\dots)$. For a set ${\mathcal{A}}$, we denote a set of all finite sequences with elements from ${\mathcal{A}}$ by ${\overline{\mathcal{A}}}={\left\{{()}\right\}}\cup\bigcup_{i=1}^{\infty}{\mathcal{A}}^{i}.$ For a vector $a=a_{1}^{n}\in{\overline{\mathcal{A}}}$ we denote by ${\mathbf{N}}(a)=n$ the vector length. We denote the concatenation (gluing) operator $\mathbf{g}$ as $\mathbf{g}(a_{1}^{n},b_{1}^{m})=(a_{1},a_{2},...,a_{n},b_{1},...,b_{m}),$ which is also generalized onto any number of input vectors. Also, define the function $\mathbf{g}_{n}:{\overline{\mathcal{A}}}^{n}\to{\overline{\mathcal{A}}}$, which glues $n$ vectors ${\overline{a}}_{i}\in{\overline{\mathcal{A}}}$ as $\mathbf{g}_{n}({\overline{a}}_{1}^{n})=\mathbf{g}({\overline{a}}_{1},{\overline{a}}_{2},...,{\overline{a}}_{n}).$ The inverse image $\mathbf{g}^{-1}_{n}(a)$ is the set of all possible ways of cutting $a\in{\overline{\mathcal{A}}}$ into $n$ subvectors, i.e., $\mathbf{g}^{-1}_{n}(a)={\left\{{{\overline{a}}_{1}^{n}|\mathbf{g}_{n}({\overline{a}}_{1}^{n})=a}\right\}}$. For a set ${\mathcal{S}}$, we denote by ${\mathbb{D}}_{\mathcal{S}}$ the set of all distributions over ${\mathcal{S}}$, i.e.
@@
The probability distribution for any probability, expectation or entropy is written as a subscript. The values that are not known are interpreted as random variables, and it is assumed that we take summation/expectation over all their possible values. For example, if we have a conditional distribution $W(u,x|v,y)$, where $u\in{\mathcal{U}},v\in{\mathcal{V}},x\in{\mathcal{X}},y\in{\mathcal{Y}}$, and the values of $u$ and $v$ have been defined previously, then
@@
We write ${\mathbf{H}}_{A}{\left[{B}\right]}={\mathbf{E}}_{A}{\left[{-\log_{2}B}\right]}$ for entropy-like expressions. If $A=B$, then we omit the subscript for simplicity:
@@
II-B Channels and channel sequences
@@
A channel sequence $V$ is defined as an infinite sequence $V^{\infty}$ of channels $V_{i}:{\mathcal{X}}^{i}\rightsquigarrow{\mathcal{Y}}_{i}$. Note that for any $i$, the output set ${\mathcal{Y}}_{i}$ is arbitrary, but the input set for the $i$-th channel is ${\mathcal{X}}^{i}$, so the $i$-th channel transmits exactly $i$ symbols from ${\mathcal{X}}$.
II-C Mutual information capacity
The mutual information density of a channel $W$ for an input distribution $p\in{\mathbb{D}}_{\mathcal{X}}$ and given values of an input $x\in{\mathcal{X}}$ and an output $y\in{\mathcal{Y}}$ is defined as:
@@ II-C Mutual information capacity
The mutual information capacity/i-capacity [4] of channel $W$ is the maximum possible mutual information of $W$ under any input distribution, and is given by [^3] In [4], the setup is for the much more general case, and “$\sup$” over the input distributions is used instead of “$\max$”. In our setup we assumed (see Section II-B) that ${\mathcal{X}}$ is finite and ${\mathcal{Y}}$ is countable, so the maximum is always achieved as a maximum of a continuous function over a compact set.
@@ II-D Coding capacity
Consider a channel sequence $V^{\infty}$. We call $R\in{\mathbb{R}}_{\geq 0}$ an achievable rate over $V$, if
@@
We denote by ${\mathbf{R}}(V)$ the set of achievable rates $R$ for channel sequence $V$. By the coding capacity/c-capacity ${\mathbf{C}}(W)$ of the channel sequence $V$ we mean the supremum of all achievable rates [5]:
${\mathbf{C}}(V)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\sup{\mathbf{R}}(V)$ (8)
The existence of a $(W,M,\varepsilon)$-coding scheme obviously implies the existence of a $(W,M^{\prime},\varepsilon^{\prime})$-coding scheme for any $M^{\prime}\leq M$ and $\varepsilon^{\prime}\geq\varepsilon$. So, if $R\in{\mathbf{R}}(V)$, then $R^{\prime}\in{\mathbf{R}}(V)$ for any rate $R^{\prime}<R$. Thus, the set ${\mathbf{R}}(V)$ is a connected set, i.e., there always exists $\alpha\in[0,\log_{2}|{\mathcal{X}}|]$, such that ${\mathbf{R}}(V)\in{\left\{{[0,\alpha],[0,\alpha)}\right\}}$. This implies that ${\mathbf{C}}(V)=\alpha$, i.e., any channel sequence has the c-capacity.
In [3], the coding capacity (c-capacity) ${\mathbf{C}}(V)$ is derived in terms of mutual information density:
${\mathbf{C}}(V)=\sup{\left\{{\alpha\left|\lim_{n\to\infty}~{}\min_{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\Pr_{p_{n}\circ V_{n}(x,y)}{\left\{{\frac{\mathbf{i}(V_{n},p_{n},x,y)}{n}\leq\alpha}\right\}}=0\right.}\right\}}$ (9)
II-E Information stability
A channel sequence $V$ is called $J$-information stable [2], where $J=J^{\infty}$ with $J_{n}\to\infty$, if there exist distributions $p^{*}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$ and sequence $\delta$ with $\delta_{n}>0$, $\delta_{n}\to 0$, such that
$\Pr_{p^{*}_{n}\circ V_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{J_{n}}-1}\right|>\delta_{n}}\right]}<\delta_{n}.$ (10)
If $J_{n}={\mathbf{I}}(V_{n})$ then $V$ is called information-stable. We denote by ${\mathbf{I}}{\mathbf{S}}(V)$ the fact that $V$ is information-stable, and, moreover, the i-capacity ${\mathbf{I}}(V)$ exists. In the case of ${\mathbf{I}}{\mathbf{S}}(V)$, under our general assumption that ${\mathcal{X}}$ is finite, (10) can be rewritten as
$\forall\delta>0:\lim_{n\to\infty}~{}\Pr_{p^{*}_{n}\circ V_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]}=0.$ (11)
Proposition 1 (Theorems 3.1 and 3.2 in [2]).
${\mathbf{I}}{\mathbf{S}}(V)\iff{\mathbf{C}}(V)={\mathbf{I}}(V)$.
For further derivations we need the following sufficient condition of information stability.
Proposition 2 (A sufficient condition for strict information stability).
If there exists ${\mathbf{I}}(V)$ and a sequence of distributions $p^{*}$, for which
$\lim_{n\to\infty}~{}{\mathbf{E}}_{p^{*}_{n}\circ V_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0,$ (12)
then ${\mathbf{I}}{\mathbf{S}}(V)$.
Proof.
Note that for any $\delta>0$,
${\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}>\delta\cdot\Pr{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]},$
so
$0\leq\lim_{n\to\infty}~{}\Pr{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]}\leq\frac{1}{\delta}\cdot\lim_{n\to\infty}~{}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0,$
which is exactly the information stability condition (11). ∎
III Functions of channels and their properties
For a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and function $f:{\mathcal{Y}}\to{\mathcal{Z}}$ we write $W^{\prime}=f(W)$ to refer to a channel $W^{\prime}:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, defined as a composition of channel $W$ and function $f$. That is,